# InferLog: Accelerating LLM Inference for Online Log Parsing via ICL-oriented Prefix Caching

### v1

@@ 3.2.2. ICL Table Management
When handling a new request, we first try to perform Prefix-based Matching to find , then update the elements in the ICL Table after processing the request. We have established the following effective rules to refine the ICL and update the ICL table in different cases:
@@
Whenever the ICL Table reaches its capacity, the eviction of the head element is carried out based on the LRU policy. Notably, InferLog offline determines the table length based on the profiling of distribution of requested tokens and remaining GPU memory which can be easily obtained in the inference system(Kwon et al., 2023).
@@ 3.3.1. Problem Formulation
is a black-box function, the value can be observed based on a complete workload replay. While increasing the sampling number is benefit for capture configuration-performance plane to identify the optimal configuration, considering the requirement for near real-time responsiveness in online tuning scenario(Sen et al., 2015; Zhang et al., 2021), the search process for sampling configurations is under budget constraints . Although many metrics exist to evaluate performance in LLM inference systems like TTFT and TPOT(Cheng et al., 2024). This paper focuses on optimizing total completion time as a single-objective optimization process because it is easy to measure and optimize.
@@ 3.3.2. Attention MAML
Here is the base learning rate of the learner. The loss function is defined as MSE between the predicted performance and the real one.
In the outer-loop update, we further calculate the expected meta-loss across all tasks according to the post-update parameters , this meta-loss from validation data is minimized to optimize the initial parameter value :
@@
Automatic Weights Learning. Inspired by prior work(Wang et al., 2024; Hao et al., 2023), InferLog implements the attention mechanism(Vaswani et al., 2017) to automatically learn optimal weight distributions, indicating the importance of meta-tasks in relation to target tasks. Specifically, it explores the potential impact relationships between meta-tasks and the target task by learning the dynamic attention among workload feature vectors , thereby facilitating task-specific model initialization for the target domain. To acquire adaptive weights within the context of MAML, we assign a separate attention module with learnable parameters , the attention weight between target task and meta task compares the workload feature vector and through a - system:
@@ 3.3.3. Enhancing SMBO with AttMAML
Sequential Model-based Optimization (SMBO) is a powerful method for configuration tuning(Feurer et al., 2015; Wang et al., 2021). The standard workflow of SMBO is: 1) Utilizing the surrogate model and the acquisition function to determine the next configuration. 2) Evaluating the newly generated configuration to obtain corresponding performance. 3) Adding the new sample into the observation dataset and then updating the surrogate model. This process is repeated until the optimal solutions are identified or the search cost exhaustion. the workflow of configuration tuning in InferLog is shown in Fig. 10.
@@ 4. Experimental Evaluation
RQ2: How does InferLog adapt to different LLMs and dynamic invocations?
@@ 4.1. Experimental Settings
Environment and Implementation. We conduct our experiments on a NVIDIA Tesla A40 48G GPU. The CUDA version on the NVIDIA platform is 12.4. The CPU is Intel(R) Xeon(R) Gold 6326 CPU @2.90GHz, 128 GiB host memory. The operation system is Ubuntu 20.04. We implement InferLog with Python 3.11. We adopt vLLM as the LLM inference system, the version is v0.6.3. We use Qwen2.5-14B(Qwen Team, 2024) as underlying LLM because it is a relatively small yet powerful model, balancing performance and efficiency in various tasks. In RQ2, we evaluate InferLog by replacing Qwen2.5 with other open-source LLMs.
During the AttMAML training process, we divided the log datasets into two groups: half for meta-training and the remaining half for testing. To obtain prior tuning knowledge, we apply the Latin hypercube sampling(LHS)(McKay et al., 2000) to uniformly sample 100 points within the configuration space under each meta-training task for training AttMAML. Besides, in training phase, the sizes of the support set and query set in MAML are 15 and 45, the meta learning rate is set to 0.0001 and the base learning rate is set to 0.001. The model consists of two fully connected layers, with 64 neurons each, and the attention matrix has a dimension of 32. We report the performance of optimal configuration within 15 steps.
@@
Configurations and Workloads. We optimize three critical scheduling-related configuration parameters in vLLM as listed in Tab.1, where #Max represents max-position-embeddings of LLM. To create prompt, we samples 200 logs from each log dataset to construct the candidate set and select 5 labeled examples from the candidate set in ICL following DivLog(Xu et al., 2024b). We conduct optimization on a fixed batch of 2000 log parsing requests processed with a fixed concurrency level of 100 by default. We use Python module to simulate concurrency online requests. In RQ2, we evaluate it’s adaptability under dynamic concurrency.
@@
Baselines. We compare InferLog with the following baselines. (1)Default: we treat vLLM(Kwon et al., 2023) with default configurations as the the basic system. (2)PC(Caching, 2024) uses default prefix caching technique for optimizing inference latency in vLLM. (3)PC-Tune: we adopt random search(Bergstra and Bengio, 2012) to optimize inference configuration parameters based on PC. (4)PC-Chunk(Agrawal et al., 2023) optimizes context filling efficiency by processing input text in chunks, enhancing generation speed and memory usage. (5)BatchLLM (Zheng et al., 2024) designs prefix-sharing group based scheduling and request groups reordering strategies to improve the prefix caching ratio. Beside, it also determines the optimal chunk size of chunked-prefills(Agrawal et al., 2023) scheduling policy to increase the throughput. Notably, except for the default, all other approaches have the prefix caching option enabled.
@@ 5.1.1. Inference Performance Improvement
Specifically, the default inference does not employ any optimization techniques, and the system fully executes the request inference process, resulting in the worst inference performance. Due to the fact that numerous log parsing requests with same log templates in log parsing task, they tend to opt for the similar demonstrations in ICL. Therefore, employing PC can lead to a significant acceleration in inference performance. However, PC(Caching, 2024), PC-Tune and PC-Chunk can only strictly reuse the KV cache with identical prefixes, which limits the reuse of KV cache between requests. PC-Chunk(Agrawal et al., 2023) utilize chunk-prefill(Agrawal et al., 2023) technique to divide requests into chunks and execute them in parallel. However, the chunking operation introduces additional overhead, which is not conducive to the execution of short requests like log parsing, so it leads to a lower throughput than PC and PC-Tune. BatchLLM(Zheng et al., 2024) performs better than other methods because it reorders the requests to improve the prefix cache hit rate. However, BatchLLM is unable to perform fine-grained ICL reordering within individual requests, leading to suboptimal KV cache reuse. Additionally, the adjustment of request order also introduces waiting delays. In contrast, InferLog adjusts the ICL examples of each incoming request, enabling optimal prefix cache utilization, and further optimizes the configuration to enhance performance.
@@ 5.2.2. Performance Under Different LLMs
To validate LLM generalization, we evaluate two representative LLMs with divergent architectures: (1) LLaMA3-8B(AI@Meta, 2024), and (2) the sparse Mixture-of-Experts (MoE)-based Mixtral-7B(Jiang et al., 2024b). As shown in Fig.13, InferLog achieves consistent performance enhancements across evaluated models. Specifically, it surpasses the baselines by 40.53% reduction on p95 latency and 1.45 throughput for LLaMA3-8B.
Due to the computational and memory access overhead introduced by the MOE mechanism, the inference speed of Mixtral model is slower than that of LLaMA3, InferLog achieves 20.52% reduction on p95 latency and 1.26 throughput for Mixtral-7B. This provides compelling evidence that InferLog is capable of effectively accelerating a variety of LLMs with diverse architectures.
@@ 5.3. Compatibility Evaluation (RQ3)
In this RQ, our primary focus is on assessing the compatibility of our method with current LLM-based log parsers. Specifically, recent studies have proposed methods to enhance the efficiency of LLM-based log parsing by analyzing the characteristics of log parsing tasks. For instance, LILAC(Jiang et al., 2024a) adopts an adaptive parsing cache to store the template to prevent duplicate queries of the LLM, while LogBatcher(Xiao et al., 2024) employs a clustering method and generate log template for each partition. Therefore, we want to see if our method can work together with the current approaches and further improve their inference efficiency. We evaluate the performance of current LLM-based log parsers: DivLog(Xu et al., 2024b), LILAC(Jiang et al., 2024a) and LogBatcher(Xiao et al., 2024) with the acceleration by the Inferlog. To avoid the additional overhead of the log parser itself, for DivLog, we query all log parsing requests in the dataset. For LogBatcher, we first offline partition logs and perform only one query on the logs of each partition. For LILAC, we only query LLM for log parsing requests with different templates. The number of examples is set to 5.
The average results across 16 datasets are illustrated in Tab.3. It is obvious that with InferLog, the performance of current log parsers has been significantly enhanced. Specifically, for DivLog, the performance improvement is the most pronounced, the reason is that a large batch of queries have brought performance bottlenecks to the LLM inference system. Through PAIR and configuration optimization by InferLog, the latency is reduced from 22.89s to 6.43s, marking a 71.90% decrease, while the throughput improvement is 301.83%. For LILAC and LogBatcher, which eliminate redundant LLM queries and focus on parsing unseen logs, the average number of log parsing queries to the LLM is reduced by 23. Nonetheless, InferLog still achieves a performance improvement of 50.17% and 49.94% in latency reduction and 66.07% and 78.31% in throughput increase for LILAC and LogBatcher, respectively. The reason is that there are still reusable KV cache in the common instruction and ICL examples among different log parsing requests, so InferLog effectively boosts prefix cache hit rates and optimizes inference performance. Overall, InferLog can improve the efficiency of LLM-based log parsers and is compatible with existing techniques that aim to reduce LLM invocation times.
| p95-Latency | Throughput | |
@@
| p95-Latency | Throughput | |
@@ 5.4.1. Contribution of PAIR
The dynamic nature of ICL in log parsing poses significant challenges for benefiting from PC techniques. PAIR is a crucial module that enables a significant part of ICL tokens to leverage the historical KV cache, thereby bypassing redundant computations cross online log parsing requests. As shown in Tab.4, when PAIR is removed, LLM can only reuse the KV cache of common instruction in user requests, due to its strict requirement for identical prefixes. Without it, the inference latency increases by 46.97% and processing throughput drops by 32.02%.
PC Hit Rate. To better understand the performance drop, Fig.14 illustrates the average PC hit rates across all datasets. Obviously, PAIR achieves the highest hit rate, with average 1.23, 1.16 and 1.18 higher than naive PC(Caching, 2024), PC-Chunk(Agrawal et al., 2023) and BatchLLM(Zheng et al., 2024), and reaching up to 2.1 on the HDFS dataset compared with PC, confirming the effectiveness of PAIR.
Impact on Parsing Accuracy. To delve deeper into the parsing accuracy, we analysis the hidden representations when implementing PAIR by Qwen2.5-14B. First, we compared the representations of input tokens where examples were positioned in various orders, which reflects the impact of Reordering. To create different orders, we adopt the same setting from prior work(Xiang et al., 2024). As depicted in Fig.16, it presents the heatmaps of average representations similarities across different example positions of Apache dataset. It is noteworthy that the cosine similarity across representations at any position typically exceeded 0.80 and up to 0.86, that is considered to possess sufficient robustness to counterbalance the demonstration order sensitivity of decoder-only LLMs according to prior work(Xiang et al., 2024). We believe that with the enhancement of LLM capabilities, their context understanding and noise resistance will improve, rendering them more robust to variations in ICL order. Furthermore, carefully designed prompt engineering also reduces the sensitivity of LLM to input order, enables LLM to comprehend tasks at a semantic level, thereby diminishing sensitivity to sequence.
Second, we compared ICL after Modifying with the original ICL on the entire prompt representation. The average similarity of output representations from the last layer is 0.9977 and 0.9983 for HealthApp and Apache dataset, respectively (omitted in the figure). In general, the semantic impact of our PAIR strategy on the prompt is not obvious, and it still maintains the high level log parsing accuracy.
@@ 5.4.2. Contribution of Configuration Tuning
Prediction Accuracy of AttMAML. InferLog innovatively integrates the attention mechanism into meta-learning framework, enabling the MAML algorithm to automatically identify which meta-training tasks are more critical for the target tasks. This enhancement not only boosts the precision of the MAML algorithm but also enhances its ability to rapidly adapt to target tasks. Specifically, we conducted meta-learning on eight tasks and evaluated the few-shot learning performance on another eight tasks. The average mean absolute percentage error (MAPE) on target tasks are shown in Fig.15, we find that (i) compared to Random Forest, the representation of machine learning, AttMAML demonstrates a 81.76% reduction on MAPE, indicating it’s ability to transfer knowledge effectively in facing new scenarios. (ii) AttMAML achieves a 32.74% MAPE reduction over the standard MAML algorithm on average, emphasizing the effectiveness of the attention mechanism in improving meta-training performance.
Online Tuning Efficiency. Efficiency is a crucial consideration in recent configuration tuning methods(Dou et al., 2022; Shen et al., 2023). To evaluate it, we compare InferLog with several prominent algorithms: LHS(McKay et al., 2000), BO with Gaussian Processes (BO-GP)(Rasmussen, 2003), DDPG(Lillicrap et al., 2015), and the SOTA method for LLM configuration tuning SCOOT(Cheng et al., 2024), the optimization was conducted within a sampling budget of 15 steps. We show tuning results of three datasets in the Fig.17. InferLog outperforms other methods by requiring 40.74%, 53.85% and 79.68% fewer iterations to determine the optimal configuration on HPC, HDFS and Linux, resulting in a performance improvement of 1.07, 1.05 and 1.13. The reason is that InferLog use a warm-start strategy to leverage historical optimal configurations from similar workloads to achieve rapid early convergence. Additionally, AttMAML-enhanced SMBO enables fine tune meta-model within few steps in new enviroments, efficiently guiding configurations tuning process.

### v3

@@ 3.2.2. ICL Table Management
When handling a new request, we first try to perform Prefix-based Matching to find , then update the elements in the ICL Table after processing the request. We have established the following effective rules to refine the demonstration set and update the ICL table in different cases:
@@
Whenever the ICL Table reaches its capacity, the eviction of the head element is carried out based on the LRU policy. Notably, InferLog offline determines the table length based on the profiling of distribution of input tokens and remaining GPU memory which can be easily obtained in the inference system(Kwon et al., 2023).
@@ 3.3.1. Problem Formulation
is a black-box function, the value can be observed based on a complete workload replay. While increasing the sampling number is benefit for capture configuration-performance plane to identify the optimal configuration, considering the requirement for near real-time responsiveness in online tuning scenario(Sen et al., 2015; Zhang et al., 2021), the search process for sampling configurations is under budget constraints . Although many metrics exist to evaluate performance in LLM inference systems like TTFT and TPOT(Cheng et al., 2024), this paper focuses on optimizing total completion time for a batch of requests as a single-objective optimization process because it is easy to measure and stable to optimize.
@@ 3.3.2. Attention MAML
Here is the base learning rate of the learner. The loss function is defined as MSE between the predicted performance and the real one. In the outer-loop update, we further calculate the expected meta-loss across all tasks according to the post-update parameters , this meta-loss from validation data is minimized to optimize the initial parameter value :
@@
Automatic Weights Learning. Inspired by prior work(Hao et al., 2023), InferLog implements the attention mechanism(Vaswani et al., 2017) to automatically learn optimal weight distributions, indicating the importance of meta-tasks in relation to target tasks. Specifically, it explores the potential impact relationships between meta-tasks and the target task by learning the dynamic attention among workload feature vectors , thereby facilitating task-specific model initialization for the target domain. To acquire adaptive weights within the context of MAML, we assign a separate attention module with learnable parameters , the attention weight between target task and meta task compares the workload feature vector and through a - system:
@@ 3.3.3. Enhancing SMBO with AttMAML
Sequential Model-based Optimization (SMBO) is a powerful method for configuration tuning(Feurer et al., 2015; Wang et al., 2021). The standard workflow of SMBO is: 1) Utilizing the surrogate model and the acquisition function to determine the next configuration. 2) Evaluating the newly generated configuration to obtain corresponding performance. 3) Adding the new sample into the observation dataset and then updating the surrogate model. This process is repeated until the optimal solutions are identified or the search cost is exhausted. The workflow of configuration tuning in InferLog is shown in Fig. 10.
@@ 4. Experimental Evaluation
RQ2: How does InferLog adapt to different LLMs and dynamic workloads?
@@ 4.1. Experimental Settings
Environment and Implementation. We conduct our experiments on NVIDIA Tesla A40 48G GPUs. The CUDA version on the NVIDIA platform is 12.4. The CPU is Intel(R) Xeon(R) Gold 6326 CPU @2.90GHz, 128 GiB host memory. The operation system is Ubuntu 20.04. We implement InferLog with Python 3.11. We adopt vLLM-v0.6.3.post1 as the LLM inference system. We use Qwen2.5-14B(Qwen Team, 2024) as underlying LLM because it is a relatively small yet powerful model, balancing performance and efficiency in various tasks. In RQ2, we evaluate InferLog by replacing Qwen2.5 with other open-source LLMs.
During the AttMAML training process, we divided the log datasets into two groups: half for meta-training and the remaining half for testing. To obtain prior tuning knowledge, we apply the Latin hypercube sampling(LHS)(McKay et al., 2000) to uniformly sample 100 points within the configuration space under each meta-training task for training AttMAML. Besides, in the training phase, the sizes of the support set and query set in MAML are 15 and 45, the meta learning rate is set to 0.0001 and the base learning rate is set to 0.001. The model consists of two fully connected layers, with 64 neurons each, and the attention matrix has a dimension of 32. We report the performance of optimal configuration within 15 steps.
@@
Configurations and Workloads. We optimize three critical scheduling-related configuration parameters in vLLM as listed in Tab.1, where #Max represents max-position-embeddings of LLMs. To create prompt, we samples 200 logs from each log dataset to construct the candidate set and select 5 labeled examples from the candidate set in ICL following DivLog(Xu et al., 2024b). We conduct optimization on a fixed batch of 2000 log parsing requests processed with a fixed concurrency level of 100 by default. We use Python module to simulate concurrency online requests. In RQ2, we evaluate it’s adaptability under dynamic concurrency.
@@
Baselines. We compare InferLog with the following baselines. (1)Default: we treat vLLM(Kwon et al., 2023) with default configurations as the the basic system. (2)PC(Caching, 2024) uses default prefix caching technique for optimizing inference latency in vLLM. (3)PC-Tune: we adopt random search(Bergstra and Bengio, 2012) to optimize inference configuration parameters based on PC. (4)PC-Chunk(Agrawal et al., 2023) optimizes context filling efficiency by processing input text in chunks, enhancing generation speed and memory usage. (5)BatchLLM (Zheng et al., 2024) designs prefix-sharing group based scheduling and request groups reordering strategies to improve the prefix caching ratio. Beside, it also determines the best chunk size of chunked-prefills(Agrawal et al., 2023) policy to achieve the optimal throughput. Notably, except for the default, all other approaches have the prefix caching option enabled.
@@ 5.1.1. Inference Performance Improvement
Specifically, the default inference does not employ any optimization techniques, and the system fully executes the request inference process, resulting in the worst inference performance. Due to the fact that numerous log parsing requests with same log templates in log parsing task, they tend to opt for the similar demonstrations in ICL. Therefore, employing PC can lead to a significant acceleration in inference performance. However, PC, PC-Tune and PC-Chunk can only strictly reuse the KV cache with identical prefixes, which limits the reuse of KV cache between requests. PC-Chunk(Agrawal et al., 2023) utilize chunk-prefill(Agrawal et al., 2023) technique to divide requests into chunks and execute them in parallel. However, the chunking operation introduces additional overhead, which is not conducive to the execution of short requests like log parsing, so it leads to a lower throughput than PC and PC-Tune. BatchLLM(Zheng et al., 2024) performs better than other methods because it reorders the requests to improve the prefix cache hit rate. However, BatchLLM do not perform fine-grained ICL reordering within individual requests, leading to suboptimal KV cache reuse. Additionally, the adjustment of request order also introduces waiting delays. In contrast, InferLog adjusts the ICL examples of each incoming request, enabling optimal prefix cache utilization, and further optimizes the configuration to enhance performance.
@@ 5.2.2. Performance Under Different LLMs
To validate LLM generalization, we evaluate two representative LLMs with divergent architectures: (1) LLaMA3-8B(AI@Meta, 2024), and (2) the sparse Mixture-of-Experts (MoE)-based Mixtral-7B(Jiang et al., 2024b). As shown in Fig.13, InferLog achieves consistent performance enhancements across evaluated models. Specifically, it surpasses the baselines by 40.53% reduction on p95 latency and 1.45 throughput for LLaMA3-8B. Also, InferLog achieves 20.52% reduction on p95 latency and 1.26 throughput for Mixtral-7B. This provides compelling evidence that InferLog is capable of effectively accelerating a variety of LLMs with diverse architectures.
@@ 5.3. Compatibility Evaluation (RQ3)
In this RQ, our primary focus is on assessing the compatibility of our method with current LLM-based log parsers. Specifically, recent studies have proposed methods to enhance the efficiency of LLM-based log parsing by analyzing the characteristics of log parsing tasks. For instance, LILAC(Jiang et al., 2024a) adopts an adaptive parsing cache to store the template to prevent duplicate queries of the LLM, while LogBatcher(Xiao et al., 2024) employs a clustering method and generate log template for each partition. Therefore, we want to see if our method can work together with the current approaches and further improve their inference efficiency. We evaluate the performance of current LLM-based log parsers: DivLog(Xu et al., 2024b), LILAC(Jiang et al., 2024a) and LogBatcher(Xiao et al., 2024) with the acceleration by the InferLog. For DivLog, we query all log parsing requests in the dataset. For LogBatcher, we first offline partition logs and perform only one query of each partition. For LILAC, we only query LLM for log parsing requests with different templates because it caches historical templates.
The average results across 16 datasets are illustrated in Tab.3. It is obvious that with InferLog, the performance of current log parsers has been significantly enhanced. Specifically, for DivLog, the performance improvement is the most pronounced, the reason is that a large batch of queries have brought performance bottlenecks to the LLM inference system. Through PAIR and configuration optimization by InferLog, the latency is reduced from 22.89s to 6.43s, marking a 71.90% decrease, while the throughput improvement is 301.83%. For LILAC and LogBatcher, they eliminate redundant LLM queries and focus on parsing unseen logs, the average number of log parsing queries to the LLM is reduced by 23. Nonetheless, InferLog still achieves a performance improvement of 50.17% and 49.94% in latency reduction and 66.07% and 78.31% in throughput increase for LILAC and LogBatcher, respectively. Overall, InferLog can improve the efficiency of LLM-based log parsers and is compatible with existing techniques that aim to reduce numbers of LLM invocation.
| p95-Latency(s) | Throughput | |
@@
5.4. Ablation Study (RQ4)
| p95-Latency(s) | Throughput | |
@@ 5.4.1. Contribution of PAIR
The dynamic nature of ICL in log parsing poses significant challenges for benefiting from PC techniques. PAIR is a crucial module that enables a significant part of ICL tokens to leverage the historical KV cache, thereby bypassing redundant computations cross online log parsing requests. As shown in Tab.4, when PAIR is removed, LLM can only reuse the KV cache of common instruction in user requests. Without it, the inference latency increases by 46.97% and processing throughput drops by 32.02%.
PC Hit Rate. To better understand the performance drop, Fig.14 illustrates the average PC hit rates across all datasets. Obviously, PAIR achieves the highest hit rate, with average 1.23, 1.16 and 1.18 higher than naive PC(Caching, 2024), PC-Chunk(Agrawal et al., 2023) and BatchLLM(Zheng et al., 2024), and reaching up to 2.1 on the HDFS dataset compared with PC(Omitted from figures), confirming the effectiveness of PAIR.
Impact on Parsing Accuracy. To delve deeper into the parsing accuracy, we analysis the hidden representations when implementing PAIR by Qwen2.5-14B. First, we compared the representations of input tokens where examples were positioned in various orders, which reflects the impact of Reordering. To create different orders, we adopt the same setting from prior work(Xiang et al., 2024). As depicted in Fig.16, it presents heatmaps of average representation similarities across different example positions of the Apache dataset. It is noteworthy that the cosine similarity across representations at any position typically exceeded 0.80 and up to 0.86, that is considered to possess sufficient robustness to counterbalance the demonstration order sensitivity of decoder-only LLMs according to prior work(Xiang et al., 2024). We believe that with the enhancement of LLMs capability and carefully designed prompt engineering, LLMs’ context understanding and noise resistance will improve, rendering them more robust to variations in ICL order. Second, we compared ICL after Modifying with the original ICL on the entire prompt representation. The average similarity of output representations from the last layer is 0.9977 and 0.9983 for HealthApp and Apache dataset, respectively (Omitted from figures). In general, the semantic impact of PAIR on the prompt is not obvious, and it still maintains the high level log parsing accuracy.
@@ 5.4.2. Contribution of Configuration Tuning
Prediction Accuracy of AttMAML. InferLog innovatively integrates the attention mechanism into meta-learning framework, enabling the MAML algorithm to automatically identify which meta-training tasks are more critical for the target tasks. This enhancement not only boosts the precision of the MAML algorithm but also enhances its ability to rapidly adapt to target tasks. Specifically, we conducted meta-learning on eight tasks and evaluated the few-shot learning performance on another eight tasks. The average mean absolute percentage error (MAPE) on target tasks are shown in Fig.15, we find that (i) compared to Random Forest(RF), the representation of machine learning, AttMAML demonstrates a 81.76% reduction on MAPE, indicating it’s ability to transfer knowledge effectively in facing new scenarios. (ii) AttMAML achieves a 32.74% MAPE reduction over the standard MAML algorithm on average, emphasizing the effectiveness of the attention mechanism in improving meta-training performance.
Online Tuning Efficiency. Efficiency is a crucial consideration in recent configuration tuning methods(Dou et al., 2022; Shen et al., 2023; Xin et al., 2022). To evaluate it, we compare InferLog with several prominent algorithms: LHS(McKay et al., 2000), BO with Gaussian Processes (BO-GP)(Rasmussen, 2003), DDPG(Lillicrap et al., 2015), and the SOTA method for LLM configuration tuning SCOOT(Cheng et al., 2024), the optimization was conducted within a sampling budget of 15 steps. We show tuning results of three datasets in the Fig.17. InferLog outperforms other methods by requiring 40.74%, 53.85% and 79.68% fewer iterations to determine the optimal configuration on HPC, HDFS and Linux, resulting in a performance improvement of 1.07, 1.05 and 1.13. The reason is that InferLog uses a warm-start strategy to leverage historical optimal configurations from similar workloads to achieve rapid early convergence. Then, AttMAML-enhanced SMBO enables the meta-model to be fine-tuned in just a few steps within new environments, efficiently guiding the configuration tuning process.